OpenStudy (anonymous):

Convert 16cm=______mm

OpenStudy (anonymous):

10mm = 1 cm

hero (hero):

16 cm = x mm 1. Get rid of the meters by dividing both sides by m: 16c = x m Change c = 1/100, m = 1/1000 so \[\frac{16}{100} = \frac{x}{1000}\] Cross multiply: 16,000 = 100x Divide both sides by 100: 160 = x

OpenStudy (anonymous):

Yeah I got that too

OpenStudy (cgreenwade2000):

People. She is asking the same type of question and not learning anything because no one is showing her how. Let her do one on her own please. Otherwise, you guys are doing someone else's homework FOR FREE

OpenStudy (anonymous):

16cm = .16m =1.6 dm=160mm answer is already given at this point so it doesnt matter

hero (hero):

I posted the correct answer to neutralize the wrong answer